OpenStudy (anonymous):

Manny and Emma are both competing in a long jump competition at their school. Manny can jump four fifths of the distance that Emma can jump. Manny can jump 40 inches. How far can Emma jump? A. 40.8 inches B. 50 inches C. 32 inches D. 39.2 inches

OpenStudy (anonymous):

manny can jump 4/5 the distance of emma. manny jumps 40 inches.\[\frac{ 4 }{ 5 }=\frac{ 40 }{ x }\] and solve for x

OpenStudy (cwrw238):

0.8 * x = 40

OpenStudy (cwrw238):

x = 400/8 = 50

OpenStudy (anonymous):

Manny gon' lose. :(

OpenStudy (anonymous):

that makes no sense to me

OpenStudy (anonymous):

@cwrw238 what?

OpenStudy (anonymous):

do you know how to cross multiply?  so 4x=5(40) 4x=200 and divide by 4